GFTODVI(1) GFTODVI(1) NAME gftodvi-makeproofsheetsfromgenericfontfiles SYNOPSIS gftodvi[-verbose][-overflow-label-offset=real][gffilename] DESCRIPTION Thegftodviprogramconvertsagenericfont(GF)fileoutputby,forexample,mf(1),toadevice independent(DVI)file(thatcanthenbetypesetusingsamesoftwarethathasalreadybeenwrittenforTeX).ThecharactersintheGFfilewillappearoneperpage,withlabels,titles,andannotationsasspecifiedinAppendixH(HardcopyProofs)ofTheMETAFONTbook. GFtoDVIusesotherfontsinadditiontothemainGFfile.A‘gray’fontisusedtotypesetthe pixelsthatactuallymakeupthecharacter.(Wewouldn’twantallthepixelstobesimplyblack, sincethenlabels,keypoints,andotherinformationwouldbelost.)A‘title’fontisusedforthe informationatthetopofthepage.A‘label’fontisusedforthelabelsonkeypointsofthefigure. A‘slant’fontisusedtotypesetdiagonallines,whichotherwisehavetobesimulatedusinghorizontalandverticalrules.Thedefaultgray,title,andlabelfontsaregray,cmr8,andcmtt10, respectively;thereisnodefaultslantfont. Tochangethedefaultfonts,youcangivespecialcommandsinyourMetafontsourcefile,oryou canchangethefontsonline.Anonlinedialogensuesifyouendthegffilenamewitha‘/’.For example, gftodviio.300gf/ Specialfontsubstitution:grayfontblack OK;anymore?grayfontarea/home/art/don/ OK;anymore?slantfont/home/fonts/slantimagen6 OK;anymore?<RET> willuse/home/art/don/blackasthe‘gray’fontand/home/fonts/slantimagen6asthe ‘slant’font(thisnameindicatesafontforlineswithslope1/6attheresolutionofanImagen printer). Thegffilenameonthecommandlinemustbecomplete.(Theprogrampromptsyouforitifyou don’tgiveit.)Becausetheresolutionispartoftheextension,itwouldnotmakesensetoaddona defaultextensionasisdonewithTeXorDVI-readingsoftware.Theoutputfilenamedefaultsto thesamerootasthegffile,withthe‘dvi’extensionadded.Forexample,theinputfileio.2602gf wouldbecomeio.dvi. OPTIONS Withoutthe-verboseoption,gftodvioperatessilently.Withit,abannerandprogressreportare printedonstdout. Theargumentto-overflow-label-offsetspecifiesthedistancefromtherightedgeofthecharacter boundingboxatwhichtheoverflowequations(ifany)aretypeset.ThevalueisgiveninTeX points.Thedefaultisalittleovertwoinches. Seetex(1)forthedetailsofspecifyingoptions. ENVIRONMENT gftodvilooksforgffilenamefirstinthecurrentdirectory.Ifitisnotpresentthere,itusesthe environmentvariableGFFONTS.Ifthatisnotset,itusesthevariableTEXFONTS.Ifthisis set,itusesthesystemdefaults: TEXFONTS:/usr/TeX/lib/tex/fonts//:., Seetex(1)forthedetailsofthesearching. FILES /usr/TeX/lib/tex/fonts/{gray.tfm,...}thedefaultfonts /usr/TeX/lib/mf/macros/{gray.mf,...}theMetafontsources(perhaps) 7Jan92 1 GFTODVI(1) GFTODVI(1) SEEALSO tex(1),mf(1) DonaldKnuth,TheMETAFONTbook DonaldKnuthetal.,METAFONTware AUTHORS DonaldKnuthwrotetheprogram.ItwaspublishedaspartoftheMETAFONTwaretechnical report,availablefromtheTeXUsersGroup.PaulRichardsportedittoUnix. 7Jan92 2